OpenStudy (anonymous):

49b^2+41b+9 Solve showing step by step please

OpenStudy (amistre64):

what do you get when you try it?

OpenStudy (amistre64):

the most basic method would be to complete the square; which is just the long hand version of the quadratic formula

OpenStudy (anonymous):

\[49b^2+41b+9\] \[ax^2+bx+c=0\] quadratic formula:\[roots = {-b \pm \sqrt{\Delta} \over 2a} \] where,\[\Delta = b^2-4ac\]

OpenStudy (amistre64):

isnt writing delta for a shortcut, and then haveing to define delta kinda defeating the purpose of writing delta to begin with ;)

OpenStudy (anonymous):

That's an expression and not an equation

OpenStudy (anonymous):

well, it depends.. you can use Delta to find the vertex so... you never know, always good to have as a side (:

OpenStudy (amistre64):

how do you use delta, the discriminate, to find the vertex?

OpenStudy (anonymous):

After you find Delta, discriminate the points of vertex should be: y=-D/4a at x=-b/2a